OpenStudy (anonymous):

The equation of a line is y = –2x + 1. What is the equation of the line that is parallel to the first line and passes through (2, 2)? A. y = –2x + 6 B. y = 2x – 2 C. y = –2x – 2 D. y = 1/2x + 1

OpenStudy (anonymous):

@phi

OpenStudy (the_fizicx99):

Y - y1=m(x-x1) So plug in those values, Btw, in y = –2x + 1 -2x is m or the slope

OpenStudy (studygurl14):

y = -2x + 1 Parallel lines have equal slopes...thus... the new line has to begin like this... y = -2x + b Plug in (2, 2) to solve for b (2) = -2(2) + b 2 = -4 + b 6 = b Plug in b into equation... and get... y = -2x + 6 Ta-da!

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

Or is it A

OpenStudy (anonymous):

So its A! Thank you

OpenStudy (the_fizicx99):

Lol you can't give out so many answer @StudyGurl14

OpenStudy (the_fizicx99):

:P

OpenStudy (anonymous):

She didnt she showed me the steps i got it

OpenStudy (the_fizicx99):

>.> the final part lol :} Just saying, you can get suspended for that

OpenStudy (anonymous):

You can get suspended for just putting the answer.. she didnt she put the steps also

OpenStudy (the_fizicx99):

She did lol, she just put the steps showing how its done, "Let the akser take pert in his question as well, this is how they learn." - Mod ... my warning. >.< Just letting ya know :p thats all.

OpenStudy (the_fizicx99):

part*

OpenStudy (anonymous):

Well im just saying she did help me and i already had the answer before.

OpenStudy (the_fizicx99):

Ok then, =.=

OpenStudy (studygurl14):

I'm confused...so I wasn't to show how to get the answer?

OpenStudy (anonymous):

I guess your not supposed to say the answer at all

OpenStudy (the_fizicx99):

@StudyGurl14 I meant you shouldn't give out answers, let them work it out as well :p

OpenStudy (studygurl14):

oh